# From Movement Primitives to Distance Fields to Dynamical Systems

### v1

@@ III-B Robot Experiments
To answer Q3, we carry out two experiments with a 7-axis Franka robot:
@@ III-B1 Disturbance Handling
In this experiment, the robot was tasked to maintain two demonstrated S-shaped and L-shaped trajectories in task space with the dimension $D=3$, as shown in Figure 7(a)-top. Disturbances are introduced by a person physically moving the robot. In the left panel, disturbances were introduced along the trajectory, and the robot remains stationary after the disturbance. In the right panel, disturbances caused displacement along the normal direction of the reference trajectory; in response, the manipulator returned to the closest point on the reference trajectory, demonstrating its adaptability to unexpected deviations. We also demonstrate the capability of disturbance handling through a pick and place task where the demonstrated trajectory is defined in joint space with the dimension $D=7$ in Figure 7(a)-bottom.
(a) Disturbance handling.
(b) Human-robot collaboration.
Figure 7: Real-world experiments.
III-B2 Human-Robot Collaboration
We demonstrate our approach in a human-robot collaboration task where a person hands an object to a robot. Demonstrations from both the human (in task space) and the robot (in joint space) are collected and encoded using quadratic splines. The total dimension for this problem is $D=3+7=10$. A RealSense D415 camera tracks the human hand and estimates the task phase using the proposed approach, allowing the robot to move accordingly. Despite disturbances from the human and vision system, the robot maintains consistent motion, anticipating the task and adjusting its position. This adaptive response improves efficiency by enabling the robot to predict and react to human movements. Figure 7(b) illustrates this interaction, showing the robot’s continuous adaptation to the user’s hand position, see accompanying video.

### v2

@@ III-B Robot Experiments
To answer Q3, we carry out three experiments with a 7-axis Franka robot:
@@ III-B1 Disturbance Handling
In this experiment, the robot was tasked to maintain two demonstrated S-shaped and L-shaped trajectories in task space with the dimension $D=3$, as shown in Figure 8-top. Disturbances are introduced by a person physically moving the robot. In the left panel, disturbances were introduced along the trajectory, and the robot remains stationary after the disturbance. In the right panel, disturbances caused displacement along the normal direction of the reference trajectory; In response, the manipulator returned to the closest point on the reference trajectory, demonstrating its adaptability to unexpected deviations, facilitated by the distance field.
III-B2 Pick and Place
In this task, the robot was required to pick up an object from the table and place it into a bowl (shown in Figure 8-middle). The demonstrated trajectory is defined in joint space with dimension $D\!=\!7$. During execution, we manually disturbed the robot by physically dragging it from the reference path. At each iteration of the control loop, the robot estimates its task phase and the closest point on the demonstrated trajectory, which defines the dynamical system to return to the desirable positions. The robot is controlled by a torque controller in a high-compliance mode, allowing us to apply external disturbances easily. IN the experiment, the gripper actions were triggered at predefined segments of the trajectory; however, this process could be integrated more dynamically into the overall framework by estimating the task phase online.
Figure 8: Real-world experiments. Top: Disturbance handling. Middle: Pick and place. Bottom: Human-robot collaboration.
III-B3 Human-Robot Collaboration
We demonstrate our approach in a dynamic human-robot handover task. A single demonstration is first recorded, encoding both the human’s hand trajectory in task space, $f_{H}(t)$, and the robot’s synchronized joint-space motion, $f_{R}(t)$, as quadratic splines. During interaction, a camera tracks the human’s current hand position, $\mathbf{x}_{H}$. Our method is used here to project this position onto the learned human path $f_{H}$ to continuously infer the task’s current phase, $t_{\text{phase}}$. This phase variable then directly drives the robot, which is commanded to the corresponding point on its own trajectory, $\mathbf{f}_{R}(t_{\text{phase}})$. This coupling creates a fluid and anticipatory interaction where the robot dynamically synchronizes its timing with the human. It robustly adapts to unpredictable movements and sensor noise, ensuring a successful and efficient handover, as the system robustly handles unpredictable human timing and noisy sensor data. This adaptive behavior is illustrated in Figure 8-bottom and in the supplementary video.